f(x) = sin x + cos x, then find the most general solution of 


f(x) = [f(pi/10)]. [.] is greatest integer function

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear

f(pi/10) =sin18 + cos18

                 =√2 sin(45+18) =√2sin63

we know sin63>sin45=1/√2 

and sin63 <1

so     1<√2sin63<2

 so [√2sin63] =1

 

now equation become 

sinx +cosx =1

cos(x-∏/4) =1/√2

cos(x-∏/4) =cos∏/4

 x-∏/4 =2n∏ ± ∏/4

 x =∏/4 +2n∏ ± ∏/4

x = 2n∏ and 2n∏ +∏/2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin